undeadknight26 (undeadknight26):

Compose an email back assisting your friend and highlight the names of the properties of exponents when you use them.

undeadknight26 (undeadknight26):

3sqrtx^3 = x (3/3) = x x^(1/3) . x^(1/3) . x^(1/3) = (1/3 + 1/3 + 1/3 = 3/3 = 1) = x^1 = x 1/x^-1 = (x^-1)/1 = x^1 = x (11^)(sqrtx^5 . x^4 . x^2) = x^(11/11) = x

undeadknight26 (undeadknight26):

@mathstudent55 @YanaSidlinskiy @sammixboo I dunno if i completed this question or not?

sammixboo (sammixboo):

O.0

OpenStudy (factor):

wht bro guess he is commin back

OpenStudy (yanasidlinskiy):

I'm comin!!:!) Gimme a sec:) Undead!!:)

sammixboo (sammixboo):

Hold on a second

OpenStudy (anonymous):

@arabpride ? ? @Georgiapeach0106 ?

OpenStudy (factor):

@sammixboo u better ask wyappt to come back

OpenStudy (yanasidlinskiy):

I *think* Not surexD!! thoughxD!! Dis is what I got:)

undeadknight26 (undeadknight26):

He is happy by himself right now is not the time for that @factor

OpenStudy (yanasidlinskiy):

*this

OpenStudy (factor):

no i guess he will be commin back @undeadknight26

OpenStudy (yanasidlinskiy):

I just took a scrnst instead of typing it all out here. It would take forever here. I've done this before by the wayxD!!:) Found it in my "Notes" foldersxD!!

OpenStudy (factor):

@undeadknight26 it was all because of @sammixboo

sammixboo (sammixboo):

It was his choice not mine sorry. Not discussing this on a question post

OpenStudy (yanasidlinskiy):

Kinda understand Undead?

undeadknight26 (undeadknight26):

Awe some banana! Thanks and yes i do!

OpenStudy (factor):

okay @sammixboo come to history

OpenStudy (yanasidlinskiy):

Welcome!!!:) Aren't u just glad I'm online?xD!!;) Lol!! <----RandomxD!!

undeadknight26 (undeadknight26):

i am @YanaSidlinskiy XD

OpenStudy (factor):

nice try

OpenStudy (factor):

i am @undeadknight26 XD

sammixboo (sammixboo):

Go to a quieter chat